OpenStudy (zyberg):

Prove that numbers 5n + 1 and 6n + 1 are coprime between each other, when n is natural number.

OpenStudy (zyberg):

That's how I would do it: if both of them aren't coprime, then their GCD would be more than 1. And their difference would be divisible by that as well: 5n + 1 = k * A 6n + 1 = k * B (6n + 1) - (5n + 1) = k(B - A) n = k(B - A) n/(B - A) = k However, I think it's not really sufficient to prove that k could only be 1... What else would I need to do?

OpenStudy (zyberg):

A and B should be natural numbers as well.

OpenStudy (zyberg):

However, it's only my guess, so if anybody has any ideas, please share them ;)

ganeshie8 (ganeshie8):

Are you really convinced that the gcd of 'a' and 'b' is same as the gcd of 'a' and 'a-b' ?

OpenStudy (kainui):

Yeah, exactly, and you can keep going with it too, so what you did was show: \[\gcd(6n+1,5n+1) = \gcd(n,5n+1)\] but you can now subtract n from 5n+1 in the same way. We can just go ahead and repeat that step 5 times to get: \[\gcd(n,5n+1) = \gcd(n,1) = 1\] At this point you're pretty much there, the greatest common divisor of 1 with any number can really only be 1. In other words relatively prime.

OpenStudy (zyberg):

Ehh, somehow I can't really understand why what @Kainui said works... Could anyone explain? ;)

OpenStudy (zyberg):

@ganeshie8 I am not convinced. I am just convinced that gcd of 6n + 1 and 5n + 1 is k.

ganeshie8 (ganeshie8):

Yes, it shouldn't be obvious if you're seeing it for the first time today It took me a couple of weeks to fully understand this result. This is also called Euclid gcd algorithm. I'd love to explain, one moment.

ganeshie8 (ganeshie8):

Maybe let's start with a simple problem : Suppose you have two logs of lengths 12 and 15. You want to cut them into pieces of equal lengths. What's the maximum possible length of each piece ?

OpenStudy (zyberg):

so, gcd(12, 15) = 3 so, cutting 4 pieces of 3 length and 5 pieces of 3 length from another log.

ganeshie8 (ganeshie8):

Yes. May i know how you figured out the GCD ?

OpenStudy (zyberg):

12 = 3 * 2^2 15 = 3 * 5

OpenStudy (zyberg):

The only number that is in both of the lengths is 3.

ganeshie8 (ganeshie8):

Nice. Suppose you didn't have a measuring tape. You need to cut them using only saw. How would you do that ?

OpenStudy (kainui):

"I am just convinced that gcd of 6n + 1 and 5n + 1 is k." So looking at your argument here: 5n + 1 = k * A 6n + 1 = k * B (6n + 1) - (5n + 1) = k*(B - A) n = k*(B - A) Let's make it look more consistent and just relabel C = (B-A) so that we have: 5n + 1 = k * A 6n + 1 = k * B n = k * C All of these numbers have the exact same gcd, which is the value k. Before we just had A and B, but now we have C which is less than A and B. So we can do this sorta tricky business. We can instead look at the gcd of the two smallest values in that list of 3 numbers, and because they all have the same gcd=k, we don't have to worry. I have also had EXTREME troubles with this too, I am pretty sure @ganeshie8 has seen me struggle many times in the past lol.

OpenStudy (kainui):

I left out the most important step, we basically can repeat the exact same process except now we choose 5n+1 and n since those are the smallest instead of using 6n+1 and 5n+1. So we are basically starting the whole process over again if that makes sense.

OpenStudy (zyberg):

I would take 15 log and place it near 12 log, 15 - 12 = 3. So, I would get a log with length of 3 and then I would cut all wood using the first 3 log as example.

ganeshie8 (ganeshie8):

Brilliant! So you do see how/why the Euclid gcd algorithm works ?

OpenStudy (zyberg):

@Kainui Oh! That clears up some things! Thanks! (Not sure to who give the medal, because both you and @ganeshie8 are big help right now ;) )

OpenStudy (zyberg):

@ganeshie8 Yes! :) It's amazing that you were able to use such a simple example. :) Thank you very much!

ganeshie8 (ganeshie8):

Awesome :) Pls medal kai. I'm on mobile atm.. I'll give you medal once I login to my computer

jhonyy9 (jhonyy9):

so i think this is really easy understandably bc. there are (5n+1) and (6n+1) but like a first step of proof we can rewriting the (6n+1) in this way ((5n+1)+1) and from the law of natural numbers the difference betwee two naturale numbers is allways 1 - bc. (6n+1) and (5n+1) has this difference of 1 so - my opinion - for every naturale numbers n from (5n+1) and (6n+1) will result coprimes @ganeshie8

OpenStudy (zyberg):

@jhonyy9 6n + 1 isn't equal to (5n+ 1)+1. (5n+ 1)+1 = 5n + 2

jhonyy9 (jhonyy9):

yes sorry i thought it (5n+n)+1)=(5n+1)+n

OpenStudy (zyberg):

And that's why your proof isn't quite right. 5n+ 1 and 6n + 1 doesn't have to be consecutive. If they were consecutive, then you would be right, but take, for example n = 2; 5n + 1 = 11 6n + 1 = 13

jhonyy9 (jhonyy9):

or (6n+1)= ((5+1)n +1) =5n+n+1 for n=k 5k+k+1= so for k=1 this is easy understandably that 5*1+1+1 is consecutive of 5*1+1 7 consec. of 6 so in this way just we use the proof of inductive for k=k+1 we get 5(k+1) +1 and (5+1)(k+1)+1 5k+6 and 5k+k +6+1 5k+6 and 6k +7 for k=1 11 and by induction not can being proven ?

jhonyy9 (jhonyy9):

or i missed something there ???